# Stochastic Kinetics of Transcription: Analysis and Computation

### v1

@@ Abstract
Stochastic gene expression is a fundamental yet complicated problem in biophysics, aiming to quantitatively describe the central dogma. Lack of thorough analysis and computationally efficient approach for fine-grained transcription models has confined relevant research to canonical Telegraph model. We study a general model of transcription and exactly solve the time-dependent chemical master equation. An approach highlighting computational efficiency is presented and analyzed for stationary distribution of mRNA molecules. Our result features generality, outstanding computational performance and explicit error estimation, unifying existing theoretical results on transcription kinetics and exhibiting applicability in practical implementation.
@@ 1 Introduction
The central dogma states that gene expression consists of transcription and translation. Transcription, synthesis of messenger RNA (mRNA) directed by genetic information encoded in DNA, is fundamental to almost all biophysical processes. Inherent stochasticity of transcription stems from stochastic nature of biochemical reactions, and becomes dominant due to low copy number of mRNA and existence of DNA as individual molecules [1]. This intrinsic fluctuation contributes prominently to the phenotypic heterogeneity [2], and is studied extensively both experimentally and mathematically [3].
@@
The canonical mathematical formulation of transcription is Telegraph model, where gene switches memorylessly between active and inactive states [4]. Telegraph model is widely used to quantitatively study the bursting kinetics [3, 5], because it admits a concise and computationally tractable expression of steady-state distribution of mRNA molecules. However, Telegraph model is an overly simplified description of transcription [6]. Substantial experimental evidence suggests that transcription is a complex process governed by intricate mechanism [7]. In particular, holding time of either state is observed to be nonexponential [5], implying the complex structure of gene states. Numerous refined mathematical models of transcription are proposed and analyzed [8, 9, 10, 11, 12], but few obtains comprehensive and pragmatic results. More precisely, most existing studies are confined to models with specific gene-state structure and computationally feasible results are restricted to low-order summary statistics.
@@
Essentially, the difficulty lies in the intractability of chemical master equation (CME) [13]. In general, CME can rarely anticipate a time-dependent exact solution and is usually studied approximately [14] or numerically [15, 16]. Among the methods, stochastic simulation algorithm (also known as Gillespie algorithm or SSA) [15] is the only exact approach. A variety of improvements of the original SSA have been devised [15]. However, in general, such stochastic simulation methods are computationally expensive and rely heavily on empirical exploration, thus may become inefficient in practical implementation. Approaches based on queueing theory are also exploited to provide theoretical results for an extensive range of stochastic gene expression problems, circumventing chemical master equation [17, 18]. But shortcomings include lack of physical interpretation and high computational cost. Even if CME is analyzed directly, we usually consider only the steady-state solution, setting the time derivatives to zero. In such case, the existence of equilibrium is presumed, which is not necessarily valid. Even if the steady-state solution is obtained, we have limited information about long time behavior of reaction kinetics since whether the system approaches the equilibrium and the rate of approaching are obscure.
@@
In this Letter, we analyze a general model of transcription with complicated gene-state architecture and state transition upon transcription. We constructively present the exact time-dependent solution to the corresponding CME via power series method for solving differential equation. It is then shown that the probability distribution of mRNA molecules can be efficiently computed at steady state based on binomial moments. Explicit error estimation for practical implementation are also provided. Our approach is generally based on CME and generating function method, involving original matrix calculus techniques. Our result features generality, outstanding computational performance and explicit error estimation.
@@
This Letter is structured as follows. We first present the main results as stated above. The detailed derivations can be found in the Supplementary Material. Then we derive two special cases from our general result: canonical Telegraph model [4] and the circumstance under renewal condition [17]. With mathematical reduction, we show the consistency of our result with previous studies. Verification of our analytical result by SSA is also performed. Finally, we briefly show how intrinsic and extrinsic noise can be collaboratively studied in our model and generalize the Heavy-Tailed Law proposed in [19].
@@ 2 Main Results
In this Letter, we mainly study a fine-grained model of transcription with complicated gene-state architecture, as illustrated in the following reaction scheme (1).
@@
$\mathcal{S}_{i}(1\leq i\leq N)$ denotes $N$ gene-states; M denotes mRNA molecule. Reactions in the first row represent transitions among different states without transcription, while reactions in the second row are transitions with production of single mRNA molecule. The reaction in the last denotes hydrolysis of mRNA. In this model, arbitrarily many states of gene are allowed and transitions occur with or without production of mRNA molecules. Considering the standard presumption for single-step biochemical reaction [20], all the transitions are supposed to be Markovian. Let $a_{i,j}(1\leq i\neq j\leq N)$ denotes the transition rate without transcription from state $i$ to $j$ and $b_{i,j}(1\leq i,j\leq N)$ the rate with transcription from state $i$ to $j$. Hydrolysis of mRNA molecules is assumed to be a first-order Poisson process independent of gene-state, with parameter $\delta$. Let $\mathbb{P}_{(i,j)}(m,t)$ denote the conditional probability that $m$ mRNA molecules are present in the system and gene happens to be in state $j$ at time $t$, assuming gene-state to be $i$ at time 00. The CME corresponding to (1) is ($\forall$ $1\leq i\leq N$ and $m\in\mathbb{N}$)
@@
By rearranging (2), we have partial differential equation
@@
$\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}d\bm{t},$
where $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots t_{k}\leq t\}\subseteq\mathbb{R}^{k}$. A detailed proof is in the Supplementary Material [21]. Similar results can be found within the field of queueing theory [22]. We note that our derivation is based on CME and differs essentially from those in queueing theory. The existence of an analytical expression of $\mathcal{G}(t,z)$ reveals why relatively elegant results can be expected by analyzing specific models through different approaches [8, 9, 10]. Though theoretically satisfactory, (4) is too complicated for practical implementation. For mathematical and computational tractability, we instead consider the binomial moment of mRNA molecules in the time limit. Note that $D:=D_{0}+D_{1}$ is a $Q$-matrix, meaning off-diagonal elements are nonnegative and row sum is all null. Assume $D$ to be irreducible throughout the Letter. For simplicity, we assume the underlying Markov chain of gene-state has converged to equilibrium, and take the invariant distribution to be $\bm{\pi}$. This assumption is reasonable if the architecture of gene-state reconstruct substantially slow compared with transcription kinetics. We now cease tracking microstate information of gene and focus on the copy number of mRNA. By summing over current states and setting the initial distribution to be $\bm{\pi}$, we define scalar binomial moments and probability distribution of mRNA molecules ($m,n\in\mathbb{N}$)
@@
where $\bm{I}_{N}$ is the identity matrix. Note $\lambda\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for $\lambda>0$, thus nonsingular by Lévy-Desplanques theorem [23]. The derivation of (5) is in the Supplementary Material [21]. The special case where $D_{1}$ is diagonal has already been studied as a queueing system $MMPP/M/\infty$ [17], and (5) agrees with existing results [24]. As demonstrated in [25], binomial moments can be more tractable than probability distribution itself in the study of reaction system.
Furthermore, we can estimate the convergence of $\{B_{m}\}_{m\in\mathbb{N}}$. Note that $B_{0}$ is always unity by definition. For a strictly row diagonally dominant matrix $C=(c_{i,j})_{N\times N}$ with $\Theta$ defined as $\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$, $\lVert C^{-1}\rVert_{\infty}$ is bounded by $\Theta^{-1}$ from above (an analogue to Theorem 4.1.2 in [26]). The following result can confirm convergence and further delineate the velocity of convergence. To specify,
@@
In particular, $\{B_{m}\}_{m\in\mathbb{N}}$ converges to zero.
Exact probability distribution of mRNA molecules $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is uniquely determined by binomial moments $\{B_{m}\}_{m\in\mathbb{N}}$ and can be given by the identity
@@
where $\binom{m}{n}$ is the combinatorial number. Substituting (5) into (7), a theoretically exact and computationally feasible approach to obtaining stationary distribution of mRNA copy number is established. Additionally, we demonstrate that series (7) converges, proving the existence of a stationary distribution of mRNA population given arbitrary parameters in (1). The convergence of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ towards zero is also guaranteed. To quantify the asymptotic behavior of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$, we have
@@
Since the upper bound converges to zero as $n$ grows, sequence $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ tends to null. Note also that this bound becomes practical only when $n$ is relatively large. As a series, evaluating the value of $\mathcal{P}_{n}$ based on (7) involves determining the truncation of terms. To ensure satisfactory accuracy, a priori estimation of truncation error is expected and is substantially favored over empirical choice. Nonetheless, based on our knowledge, none of existing approaches to analyzing stochastic gene expression model effectively achieve this objective. In contrast, in our model, we are able to give an explicit upper bound for truncation error. Let $\widehat{\mathcal{P}_{n}}$ denote the computed value given the truncation point $M$. Assume $M>\max\{2n+2,8\mathrm{e}\left(\lVert D_{1}\rVert_{\infty}/\delta\right)^{2}\}$. Then
$\mid\mathcal{P}_{n}-\widehat{\mathcal{P}_{n}}\mid\leq\frac{1}{\sqrt{2\pi\mathrm{e}}n!2^{M-1}}.$ (9)
@@
Conventionally, when analyzing biochemical reaction systems like (1), we assume in advance the existence of equilibrium, and calculate the steady-state solution with all the time derivatives in the CME set to zero. In this Letter, instead, we present directly an analytical formulation of the time-dependent solution and then take the limit with respect to time. Thereby, our results additionally illustrates that the reaction system (1) possesses a unique steady state and always converges to equilibrium after a sufficiently long period of time. The rate of approaching the equilibrium can also be obtained by studying the time-dependent solution (4). Moreover, explicit upper bounds (6)(8)(9) can be exploited to design a predictable computational framework for calculating the summary statistics of arbitrarily high order or the probability distribution of mRNA copy number.
@@
Figure 1: Upper Bound of Binomial Moments and Probability Distribution of mRNA molecules: In the left illustration, binomial moments and upper bound are computed according to (5) and (6), respectively. In the right illustration, the upper bound of probability distribution is given by (8). Without loss of generality, we normalize all the parameters by dividing $\delta$. Hence $\delta$ is set as unity in the following examples. In the above illustrations, parameters are both set as $D_{0}=$ $\begin{pmatrix}-7&2&3&0\\ 3&-12&1&5\\ 1&2&-12&6\\ 8&8&0&-19\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&1&0&0\\ 1&0&1&1\\ 0&1&1&1\\ 1&1&0&1\end{pmatrix}$
3 Verification by Simulation and Connection with some Known Results
3.1 Stochastic Simulation
@@
Stochastic simulation algorithm serves as a universal and exact approach to studying chemical reactions. We first verify our analytical result (5)(7) with stochastic simulation [27] of the reaction system (1). As Figure 2 demonstrates, our theoretical result of steady-state distribution of mRNA molecules agrees well with output of simulations.
@@
Figure 2: Probability Distribution of mRNA Molecules Obtained by Stochastic Simulation and Our Theoretical Result: To perform SSA, we implement Python package GillesPy2 [27]. In Example $1$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-16.01\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&15\end{pmatrix}$ The histogram is generated based on $50000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results presented in the Letter. In Example $2$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.2&0.1&0&0&0.1\\ 0.1&-2.7&0.5&0&0.1\\ 0.1&0.2&-5.4&0&0.1\\ 0.1&0.1&0.3&-3.5&0\\ 0.1&0.1&0.1&0.1&-15.4\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}0&0&1&0&1\\ 1&1&0&0&0\\ 1&2&1&1&0\\ 1&0&0&1&1\\ 0&0&0&0&15\end{pmatrix}$ The histogram is generated based on $100000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results. Each trajectory is truncated at dimensionless time $t=100$.
@@
Additionally, our model (1) possesses a significant degree of generality, covering many previous results. In this Letter, we focus on only two exemplars: the classic Telegraph model [4] and the special case where transcription becomes a renewal process [17].

### v2

# Stochastic Kinetics of mRNA Molecules in a General Transcription Model

@@ Abstract
Stochastic gene expression is a fundamental yet complicated problem in biophysics, aiming to quantitatively describe the central dogma. Lack of thorough analysis and computationally efficient approach for fine-grained transcription models has confined relevant research to canonical Telegraph model. We study a general model of transcription and exactly solve the time-dependent chemical master equation. An approach highlighting computational efficiency is presented and analyzed for stationary distribution of mRNA molecules. Distribution of mRNA copy number is found to be upper-bounded by a constant multiple of Poisson distribution. Our result features generality, profundity and outstanding computational performance, unifying existing theoretical results on transcription kinetics and exhibiting broader applicability in practical implementation.
@@ 1 Introduction
The central dogma states that gene expression consists of transcription and translation. Transcription, synthesis of messenger RNA (mRNA) directed by genetic information encoded in DNA, is fundamental to almost all biophysical processes. Inherent stochasticity of transcription stems from stochastic nature of biochemical reactions, and becomes dominant due to low copy number of mRNA and existence of DNA as individual molecules [11, 21]. This intrinsic fluctuation contributes prominently to the phenotypic heterogeneity [32], and is studied extensively both experimentally and mathematically [3]. In particular, an overarching theoretical perspective is provided in [29].
@@
The canonical mathematical formulation of transcription is Telegraph model, where gene switches memorylessly between active and inactive states [30]. Telegraph model is widely used to quantitatively study the bursting kinetics [26, 36] and transcriptional regulation [31], because it admits a concise and computationally tractable expression of steady-state distribution of mRNA molecules. However, Telegraph model is an overly simplified description of transcription [18]. Substantial experimental evidence suggests that transcription is a complex process governed by intricate mechanism [5]. In particular, holding time of either state is observed to be nonexponential [36], implying the complex structure of gene states. Numerous refined mathematical models of transcription are proposed and analyzed [4, 13, 14, 16, 20, 35, 38], but few obtains comprehensive and pragmatic results. More precisely, most existing studies are confined to models with specific gene-state structure and computationally feasible results are restricted to low-order summary statistics.
@@
Essentially, the difficulty lies in the intractability of chemical master equation (master equation or CME) [19]. In general, CME can rarely anticipate a time-dependent exact solution [9] and is usually studied approximately [2, 33] or numerically [7, 12, 25]. Among the methods, stochastic simulation algorithm (also known as Gillespie algorithm or SSA) [7] is the only exact approach. A variety of improvements of the original SSA have been devised [7]. However, in general, such stochastic simulation methods are computationally expensive and rely heavily on empirical exploration, thus may become inefficient in practical implementation. Approaches based on queueing theory are also exploited to provide theoretical results for an extensive range of stochastic gene expression problems, circumventing chemical master equation [17, 37]. But shortcomings include lack of physical interpretation and high computational cost. Even if CME is analyzed directly, we usually consider only the steady-state solution, setting the time derivatives to zero. In such case, the existence of equilibrium is presumed, which is not necessarily valid. Even if the steady-state solution is obtained, we have limited information about long time behavior of reaction kinetics since whether the system approaches the equilibrium and the rate of approaching are obscure. However, we demonstrate in this article that the chemical master equation of a general transcription model is exactly solvable.
@@
In this article, we analyze a general model of transcription with complicated gene-state architecture and state transition upon transcription. We constructively present the exact time-dependent solution to the corresponding CME via power series method for solving differential equation. It is then shown that the probability density function of mRNA molecules can be effectively investigated at steady state based on binomial moments. Sharp upper bounds for binomial moments are given based on an analogue to Theorem 4.1.2 in [8]. Subsequently, an upper bound for probability density function is also derived, which shows that probability density function of mRNA molecules is always bounded by a constant multiple of Poisson distribution. This observation is later used to generalize the Heavy-Tailed Law proposed in [10]. Explicit error estimation for practical implementation is also provided. Our approach is generally based on CME and generating function method, involving original matrix calculus techniques. Our result features generality, profundity and outstanding computational performance.
@@
This article is structured as follows. We first present the main results as stated above. The detailed derivations can be found in the Appendix. Then we derive two special cases from our general result: canonical Telegraph model [30] and the circumstance under renewal condition [37]. With mathematical reduction, we show the consistency of our result with previous studies. Verification of our analytical result by SSA is also performed. Finally, we briefly show how intrinsic and extrinsic noise can be collaboratively studied in our model and prove the generalized Heavy-Tailed Law.
@@ 2 Main Results
In this article, we mainly study a fine-grained model of transcription with complicated gene-state architecture, as illustrated in the following reaction scheme (1).
@@
$\mathcal{S}_{i}(1\leq i\leq N)$ denotes $N$ gene-states; M denotes mRNA molecule. Reactions in the first row represent transitions among different states without transcription, while reactions in the second row are transitions with production of single mRNA molecule. The reaction in the last denotes hydrolysis of mRNA. In this model, arbitrarily many states of gene are allowed and transitions occur with or without production of mRNA molecules. If necessary, enlarge $N$ to secure that each transition represents one single-step biochemical reaction. Note that reactants not directly involved in the above process have been omitted to underscore the reaction structure. For instance, if translation of protein from mRNA and degradation of protein are considered [34], the marginal distribution of mRNA molecules will be exactly our result presented in this article. Considering the standard presumption for single-step biochemical reaction [6], all the transitions are supposed to be Markovian. Let $a_{i,j}(1\leq i\neq j\leq N)$ denote the transition rate without transcription from state $\mathcal{S}_{i}$ to $\mathcal{S}_{j}$ and $b_{i,j}(1\leq i,j\leq N)$ the rate with transcription from state $\mathcal{S}_{i}$ to $\mathcal{S}_{j}$. Hydrolysis of mRNA molecules is assumed to be a first-order Poisson process independent of gene-state, with parameter $\delta$. Let $\mathbb{P}_{(i,j)}(m,t)$ denote the conditional probability that $m$ mRNA molecules are present in the system and gene happens to be in state $\mathcal{S}_{j}$ at time $t$, assuming gene-state to be $\mathcal{S}_{i}$ at time 00. According to (1), number of mRNA molecules together with gene state is a continuous-time Markov chain with state space ×$\mathbb{N}\times\{1,2,\cdots,N\}$. The corresponding CME is ($\forall$ $1\leq i,j\leq N$ and $m\in\mathbb{N}$)
@@
By rearranging (2), we have partial differential equation system
@@
$\mathrm{e}^{D_{0}t_{1}}D_{1}\mathrm{e}^{D_{0}(t_{2}-t_{1})}D_{1}\cdots\mathrm{e}^{D_{0}(t_{k}-t_{k-1})}D_{1}\mathrm{e}^{D_{0}(t-t_{k})}\mathrm{d}\bm{t},$
where $\Omega_{k}:=\{(t_{1},t_{2},\cdots,t_{k})\mid 0\leq t_{1}\leq t_{2}\leq\cdots\leq t_{k}\leq t\}\subseteq\mathbb{R}^{k}$. A detailed proof is in Appendix A. Similar results can be found within the field of queueing theory [23], which provides physical interpretation of the expression (4). We note that our derivation is based on CME and differs essentially from those in queueing theory. The existence of an analytical expression of $\mathcal{G}(t,z)$ reveals why relatively elegant results about probability density function can be expected by analyzing specific models through different approaches [4, 13, 16, 38]. Though theoretically satisfactory, (4) is too complicated for practical implementation. For mathematical and computational tractability, we instead consider the binomial moment of mRNA molecules in the time limit. Note that $D:=D_{0}+D_{1}$ is a $Q$-matrix, meaning off-diagonal elements are nonnegative and row sum is null. Assume $D$ to be irreducible throughout the article. For simplicity, we assume the underlying Markov chain of gene-state has converged to equilibrium, and take the invariant distribution to be $\bm{\pi}$. This assumption is reasonable if the architecture of gene-state reconstruct substantially slow compared with transcription kinetics. We now cease tracking microstates of gene and focus on the coarse-grained copy number of mRNA. By summing over current states and setting the initial distribution to be $\bm{\pi}$, we define scalar binomial moments and probability density function of mRNA molecules ($m,n\in\mathbb{N}$)
@@
where $\bm{I}_{N}$ is the identity matrix. Note $\lambda\bm{I}_{N}-D$ is a strictly diagonally dominant matrix for $\lambda>0$, thus nonsingular by Lévy-Desplanques theorem [15]. The derivation of (5) is in Appendix B. The special case where $D_{1}$ is diagonal has already been studied as a queueing system $MMPP/M/\infty$ [37], and (5) agrees with existing results [27]. As demonstrated in [1], binomial moments can be more tractable than probability density function itself in the study of reaction system.
Furthermore, we can estimate the long-time behavior of $\{B_{m}\}_{m\in\mathbb{N}}$. Note that $B_{0}$ is always unity by definition. For a strictly row diagonally dominant matrix $C=(c_{i,j})_{N\times N}$ with $\Theta$ defined as $\min_{1\leq i\leq N}\left(\mid c_{i,i}\mid-\sum_{j\neq i}\mid c_{i,j}\mid\right)$, $\lVert C^{-1}\rVert_{\infty}$ is bounded by $\Theta^{-1}$ from above (an analogue to Theorem 4.1.2 in Ref. [8]). The following result confirms convergence and further delineates the velocity of convergence. To specify,
@@
In particular, $\{B_{m}\}_{m\in\mathbb{N}}$ converges to zero. See the left panel in Figure 1. Proof of (6) is in Appendix C. Exact probability density function of mRNA molecules $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ is uniquely determined by binomial moments $\{B_{m}\}_{m\in\mathbb{N}}$ and can be given according to the identity
@@
where $\binom{m}{n}$ is the combinatorial number. Substituting (5) into (7), a theoretically exact and computationally feasible approach to obtaining stationary distribution of mRNA copy number is established. Additionally, we demonstrate that series (7) converges, proving the existence of a stationary distribution of mRNA population given arbitrary parameters in (1). To quantify the asymptotic behavior of $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$, we have
@@
This inequality demonstrates that probability density function of mRNA copy number is always bounded from above by a constant multiple of the density function of one Poisson distribution. To emphasize the usefulness of this upper bound, we utilize (8) to prove generalized Heavy-Tailed Law in section 4. Since the upper bound converges to zero as $n$ grows, sequence $\{\mathcal{P}_{n}\}_{n\in\mathbb{N}}$ tends to null. Note also that this bound becomes practical only when $n$ is relatively large. See the right illustration in Figure 1. Proof of (8) is in Appendix D. As a series, evaluating the value of $\mathcal{P}_{n}$ based on (7) involves determining the truncation of terms. To obtain satisfactory accuracy, a priori estimation of truncation error is expected and is substantially favored over empirical choice. Nonetheless, based on our knowledge, none of existing approaches to analyzing stochastic gene expression model effectively achieve this objective. In contrast, in our model, we are able to give an explicit upper bound for truncation error. Let $\widehat{\mathcal{P}_{n}}$ denote the computed value given the truncation point $M$. Assume $M>\max\{2n+2,8\mathrm{e}\left(\lVert D_{1}\rVert_{\infty}/\delta\right)^{2}\}$. Then
$\mid\mathcal{P}_{n}-\widehat{\mathcal{P}_{n}}\mid\leq\frac{1}{\sqrt{2\pi\mathrm{e}}n!2^{M-1}},$ (9)
which is proved in Appendix E. However, we note that, apart from truncation errors, floating-point errors can become extremely severe when reconstructing probability density function based on (7). Essentially the same issue has been observed and analyzed in [13].
@@
Conventionally, when analyzing biochemical reaction systems like (1), we assume in advance the existence of equilibrium, and calculate the steady-state solution with all the time derivatives in the CME set to zero. In this article, instead, we construct directly an analytical formulation of the time-dependent solution and then take the limit with respect to time. Thereby, our results additionally illustrates that the reaction system (1) possesses a unique steady state and always converges to equilibrium after a sufficiently long period of time. The rate of approaching the equilibrium can also be obtained by estimating the time-dependent solution (4). Moreover, explicit upper bounds (6)(8)(9) can be exploited to design a predictable computational framework for calculating the summary statistics of arbitrarily high order or the probability density function of mRNA copy number.
@@
Figure 1: Upper Bound of Binomial Moments and Probability Distribution of mRNA molecules: In the left panel, binomial moments and upper bound are computed according to (5) and (6), respectively. In the right panel, the upper bound of probability density function is given by (8). Without loss of generality, we normalize all the parameters by dividing $\delta$. Hence $\delta$ is set as unity in this and following examples. In the above illustrations, parameters are both set as $D_{0}=$ $\begin{pmatrix}-7&2&3&0\\ 3&-12&1&5\\ 1&2&-12&6\\ 8&8&0&-19\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&1&0&0\\ 1&0&1&1\\ 0&1&1&1\\ 1&1&0&1\end{pmatrix}$
3 Stochastic Simulation and Consistency with Known Results
@@
Stochastic simulation algorithm serves as a universal and exact approach to studying chemical reactions. We first verify our analytical result (5)(7) with stochastic simulation of the reaction system (1). As Figure 2 demonstrates, our theoretical result of steady-state distribution of mRNA molecules agrees well with output of simulations.
@@
Figure 2: Probability Distribution of mRNA Molecules Obtained through Stochastic Simulation and Our Theoretical Result: To perform SSA, we implement Python package GillesPy2 [24]. In Example $1$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.02&0.01&0.01\\ 0.1&-7.2&0.1\\ 0&0.01&-16.01\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}1&0&1\\ 1&5&1\\ 0&1&15\end{pmatrix}$ The histogram is generated based on $50000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results presented in the article. In Example $2$, parameters are specified as $D_{0}=$ $\begin{pmatrix}-2.2&0.1&0&0&0.1\\ 0.1&-2.7&0.5&0&0.1\\ 0.1&0.2&-5.4&0&0.1\\ 0.1&0.1&0.3&-3.5&0\\ 0.1&0.1&0.1&0.1&-15.4\end{pmatrix}$ and $D_{1}=$ $\begin{pmatrix}0&0&1&0&1\\ 1&1&0&0&0\\ 1&2&1&1&0\\ 1&0&0&1&1\\ 0&0&0&0&15\end{pmatrix}$ The histogram is generated based on $100000$ trajectories of SSA and the steam plot is illustrated according to our theoretical results. Each trajectory is truncated at dimensionless time $t=100$.
@@
Additionally, our model (1) possesses a significant degree of generality, covering many previous results. In this article, we focus on only two exemplars: the classic Telegraph model [30] and the special case where transcription becomes a renewal process [37].